OpenStudy (goformit100):

The volume of a cube is increasing at the rate of 8 cm 3 /s. How fast is the surface area increasing when the length of an edge is 12 cm?

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

Same as here http://openstudy.com/study#/updates/517df4ebe4b0be6b54ab65e2 except \[\huge \frac{dV}{dt}=9\] and x = 12

terenzreignz (terenzreignz):

\huge \frac{dS}{dt}=\frac{dS}{dV}\cdot\frac{dV}{dt}=\frac{\frac{dS}{dx}\times \frac{dV}{dt}}{\frac{dV}{dx}}

terenzreignz (terenzreignz):

Sorry... \[\huge \frac{dS}{dt}=\frac{dS}{dV}\cdot\frac{dV}{dt}=\frac{\frac{dS}{dx}\times \frac{dV}{dt}}{\frac{dV}{dx}}\]

OpenStudy (goformit100):

Thank you SIr